OpenStudy (anonymous):

evaluate the expression 3x^2+y, if x=1 and y=-3

OpenStudy (austinl):

\(3x^2+y,~x=1,~y=-3\) \(3(1)^2+(-3)=?\)

OpenStudy (anonymous):

0 ?

OpenStudy (austinl):

Yep.

OpenStudy (anonymous):

thank you

OpenStudy (austinl):

You're welcome.

OpenStudy (anonymous):

\[the expression 5\sqrt{3} - \sqrt{27} is equivalent \to ... a) 8\sqrt{3} b)-8\sqrt{3} c-2\sqrt{3} or d)2\sqrt{3} \]

OpenStudy (austinl):

What?

OpenStudy (anonymous):

the expression 5 radical 3 minus radical 27 is equivalent to ?

OpenStudy (anonymous):

a) 8 radical 3 b) -8 radical 3 c) -2 radical 3 d) 2 radical 3

OpenStudy (anonymous):

which one?

OpenStudy (austinl):

\(\sqrt{27}~\text{can be written as}~ \sqrt{9}\times\sqrt{3}\) Do you think you can solve it from there?

OpenStudy (anonymous):

I cant see what you just typed

OpenStudy (austinl):

\(\sqrt{27}~\text{can be written as}~\sqrt{9}\times\sqrt{3}\) Do you think you can solve it from there?

OpenStudy (anonymous):

is the answer D ?

OpenStudy (anonymous):

2 radical 3

OpenStudy (anonymous):

?

OpenStudy (austinl):

Yep

OpenStudy (anonymous):

yayy